40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Aug 27 09:24:17 2012, Steve, Update, SUS, EQ damage eqDamage.png
    Reply  Mon Aug 27 10:49:03 2012, Koji, Update, SUS, EQ damage 
       Reply  Mon Aug 27 15:49:46 2012, Jenne, Update, SUS, EQ damage 
Message ID: 7286     Entry time: Mon Aug 27 15:49:46 2012     In reply to: 7283
Author: Jenne 
Type: Update 
Category: SUS 
Subject: EQ damage 

Quote:

After shaking ITMX by the alignment bias in yaw, it came back.

As ETMX seems to be largely misaligned yaw (and did not come back with the alignment impact),
the condition of the magnets are not clear. Only the side OSEM is responding nicely.

Quote:

  It looks like we may lost 1 (or 3 )  magnets? Do not panic, it's not for sure

 

I tried to take some photos through the window of ETMX's chamber, to see if I could see any magnets.  What we have learned is that Jenne is still not the world's best photographer.  I was holding the camera at ~max zoom inside the beam tube between the table and the window, so that's my excuse for the photos being fuzzy.  The only thing that I can really conclude is that the magnets look like they are still there, but Jamie thinks they may be stuck on the PDs/LEDs (now looking at the photos myself, I agree, especially with UL and LR). 

It looks like the best thing to do at this point, since Koji already tried jerking ETMX around in yaw a little bit, is just wait and open the door, to see what's going on in there.  I posted the photos on Picasa:

https://picasaweb.google.com/foteee/ETMX_MaybeStuck_ThroughWindow_27Aug2012

I propose that, if the magnets are broken, we pull the ETM out of the camber and fix it up in the cleanroom while we pump back down.  This would restrict us from doing any Xarm work, but will force me to focus on DRMI, and we can put the ETM back when we vent to install the tip tilts.

ELOG V3.1.3-